OpenStudy (vheah):

How do we prove that this is a vector space? The set of real numbers R+ Addition: x+y = xy Scalar: rx = x^r

OpenStudy (518nad):

hey so you wanna show that the set of real numbers is a vector space? or real numbers applied to these rules is a vector spae?

OpenStudy (518nad):

a vector space is when, any 2 given 'vectors' under your scalar rule will produce another 'vector' in your 'vector' space

OpenStudy (vheah):

@518nad yeah i wanna show that the set of real numbers is a vector space. I'm a little confused by the concept to say "yes these set of numbers is a vector space". It's a little tricky for me to understand

OpenStudy (518nad):

okay so lets see if your set of real numbers first satisfy your scalar rule

OpenStudy (518nad):

we have to show that any possible real number we produce with the scalar rule will definitely satisfy this scalar rule

OpenStudy (518nad):

start with 1=r, x=2 1*2=2^1 true now 2*2=2^2 4=4 okay since we saw that 4 is in this vector space, that means there must exist another number like r*4=4^r in this case r=1 is the only solution

OpenStudy (518nad):

actually i dont know if that really is the only solution. there might be more let me see

OpenStudy (518nad):

4^r-4r=0 there are infinite solutions to this equations i believe, and you need to show that every one of these solutions exist

OpenStudy (518nad):

okay lets start with the general case then

OpenStudy (518nad):

oh no only 2 solutions, since this has to be a real solutions only

OpenStudy (helpwithk12):

hi

OpenStudy (vheah):

okay @518nad i remember my professor mentioning 1 being one of the solutions. It's nice to see it this way. He never rly substituted actual numbers to emphasize the equation.

OpenStudy (helpwithk12):

whats the question

OpenStudy (518nad):

it's at the top @HelpWithK12

OpenStudy (helpwithk12):

ok

OpenStudy (518nad):

pretty cool question

OpenStudy (helpwithk12):

yeah quit itrsting but i can do it

OpenStudy (helpwithk12):

hmmmmm

OpenStudy (518nad):

x+y=xy